OpenStudy (anonymous):

-4/0.8 = 2/s-1 show work please

OpenStudy (anonymous):

You want us to solve for s or simplify?

OpenStudy (anonymous):

solve for s

OpenStudy (anonymous):

-4/0.8=2/s-1 -5=2/s-1 -5(s-1)=2/s-1 X (s-1) -5s+5=2 -5s+5-5=2-5 -5s=-3 s=-3/-5 s=0.6

OpenStudy (texaschic101):

-4/0.8 = 2/(s-1) Is this your problem ? or : -4/0.8 = (2/s) - 1 or this one ?

OpenStudy (anonymous):

first one

OpenStudy (texaschic101):

dottg is correct.... -